A tank contain 500kg of liquid whose specific gravity is 2. Determine the volume of the liquid in the tank

1
Expert's answer
2022-01-31T17:30:05-0500

"Density=specific\ gravity\times{density\ of\ water}=2\times1000=2000\frac{kg}{m^3}"


"V=\frac{m}{\rho}=\frac{500\ kg}{2000\ kg/m^3}=0.25\ m^3=250\ L"


Answer: 250 L